The pendulum of certain clock has time period 2.04 what speed of run of clock? 
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
The pendulum of the clock has time period 2.04sec. 
Now, No. or oscillation in 1 day = (24 × 3600)/2 = 43200 
But, in each oscillation it is slower by (2.04 – 2.00) = 0.04sec. 
So, in one day it is slower by, 
= 43200 × (0.04) = 12 sec = 28.8 min 
So, the clock runs 28.8 minutes slower in one day.